Betty starts working on Jan 1 2019. She is able to save $4,000 per year for retirement for the first 4 years. After that she starts to save for a house, wedding, and expenses related to a family. Over the next 40 years she does not save for her retirement. If she makes on average 7% annual interest, how much will she have from this investment when she retires?

I believe the ans is: $265942.96,
Answer #1

ANSWER:

I = 7%

Savings in 1st 4 years = $4,000

future value = ?

fv = savings in 1st year(f/p,i,n) + savings in 1st year(f/p,i,n) + savings in 1st year(f/p,i,n) + savings in 1st year(f/p,i,n)

fv = 4,000(f/p,7%,43) + 4,000(f/p,7%,42) + 4,000(f/p,7%,41) + 4,000(f/p,7%,40)

fv = 4,000 * 1.8344 + 4,000 * 1.7144 + 4,000 * 1.6023 + 4,000 * 1.4974

fv = 73,377.42 + 68,577.03 + 64,090.68 + 59,897.83

fv = 265,942.96

so the savings for her are $265,942.96 when she retires.